# SlipNet: Slip Cost Map for Autonomous Navigation on Heterogeneous Deformable Terrains

### v1

@@ IV Conclusions
In this work, we developed a framework for synthetic data generation for wheel-slip prediction in deformable terrain environments using Vortex Studio. We also proposed SlipNet, which introduces adaptive reclassification by dynamically resegmenting and reclassifying unseen terrain during deployment and maintaining a history of the most relevant segmentation class. Our method outperforms the state-of-the-art, TerrainNet method for slip prediction in all 5 test case scenarios with varying soil properties.
@@
We extend our sincere gratitude to the entire team at the Khalifa University Center for Autonomous Robotic Systems (KUCARS) and Advance Research and Innovation Center (ARIC) for their invaluable support and collaboration throughout this research.

### v2

@@ IV Conclusions
In this study, we presented SlipNet, a model for predicting wheel slip in deformable terrain environments without the need for prior terrain classification. Our method employs adaptive reclassification, allowing for dynamic segmentation and classification of previously unseen terrains during deployment. Through extensive simulations using datasets generated by our high-fidelity synthetic data framework based on Vortex Studio, we demonstrated that SlipNet significantly outperforms the state-of-the-art TerrainNet method across all five test scenarios. These results underscore the effectiveness of SlipNet in handling varying soil properties and enhancing rover navigation capabilities in uncertain terrains. Future research will focus on integrating the slip cost map generated by SlipNet into learning-based autonomous navigation systems to enable real-time decision-making in complex environments.
@@
We extend our sincere gratitude to the entire team at the Khalifa University Center for Autonomous Robotic Systems (KUCARS) and Advanced Research and Innovation Center (ARIC) for their invaluable support and collaboration throughout this research. Our gratitude also extends to Khalifa University of Science and Technology for providing the essential funding and resources that were crucial for the successful completion of this work.